Mary uses a secret code to write messages. First she translates each letter of the alphabet into a number based on its position in the alphabet, so that A = 1, B = 2, C = 3, and so on. Then she multiplies this number by her favorite number and adds her mother's favorite number. Using her secret code, F becomes 41 and T becomes 83.
Find what number M becomes.
What are Mary's favorite number and her mother's favorite number?

Answer by htmentor(1343)

Let x = Mary's favorite number
Let y = Mom's favorite number
The general formula for the encoded number is x*n + y, where n is the number's
position in the alphabet
41 = 6x + y
83 = 20x + y
Subtract the 1st from the 2nd:
42 = 14x -> x = 3
Thus y = 41 - 6*3 = 23
Therefore M = 13*3 + 23 = 62